OpenStudy (anonymous):

Five candidates - A, B, C, D and E were in an election. Candidate A was elected with 30 votes, B placed second and E was in last with 3 votes. if 59 votes were cast, and no two candidates had the same number of votes, what is the smallest number of votes that B could have received ? Please show full work

OpenStudy (anonymous):

59-3-30=26 26 votes are left for B C and D (a) +(a-1)+(a-2)=26 B C D 10+9+7=27 10+9+7

OpenStudy (anonymous):

so 10

OpenStudy (anonymous):

what are you subbing for a ?

OpenStudy (anonymous):

A=30 B=10 C= D= E=3

OpenStudy (anonymous):

ok hold on im sorry i don't understand this part (a) +(a-1)+(a-2)=26 what exactly are you doing there ?

OpenStudy (anonymous):

26 is remaing votes after candidate a , b and e

OpenStudy (anonymous):

ohh ok thank you ! :D

OpenStudy (anonymous):

but wouldnt B be 7 in that case ? since it's the least amount of votes b could have received ?

OpenStudy (anonymous):

but B is highest

OpenStudy (anonymous):

** higher than C D E

OpenStudy (anonymous):

ohh true okay then thanks :D can u do me a favor and answer some of the questions i posted up http://openstudy.com/#/updates/4edec488e4b05ed8401abf67